OpenStudy (anonymous):

Find the percent of increase from 320 to 380. Round to the nearest tenth of a percent if necessary. (1 point) 18.8% 84.2% 15.8% 0.2%

jimthompson5910 (jim_thompson5910):

Similar Problem (Same idea, the numbers have changed) Find the percent of increase from 350 to 390. Round to the nearest tenth of a percent if necessary. Percent Increase = [(New - Old)/Old] * 100 Percent Increase = [(390-350)/350] * 100 Percent Increase = [40/350] * 100 Percent Increase = 0.114285714285714 * 100 Percent Increase = 11.4285714285714% Percent Increase = 11.4% (round to the nearest tenth) Note: this is NOT the answer to your problem, it's just an example

OpenStudy (anonymous):

i need the answer